# Larger Language Models Don’t Care How You Think: Why Chain-of-Thought Prompting Fails in Subjective Tasks thanks: Funded in part by DARPA under contract HR001121C0168

### v1

@@ V Conclusion
In this work, we evaluate whether CoT displays the same priors as simple ICL. We find, surprisingly, that larger language models do have reasoning priors that ossify their generated reasoning and posterior predictions, virtually disregarding the evidence in the form of reasoning and labels provided in the prompt. Given that the performance of LLMs is inferior to traditional methods in the tasks we study, this is unlikely to be an artifact of data memorization. We conclude, therefore, that reasoning in the form of CoT prompting is unable to overcome the pull of the priors in larger language models, leading to suboptimal performance on subjective tasks, in contrast to other tasks where labels, and in turn reasoning, remain more consistent across datasets, the validity of a label can be asserted more formally and consistently, and therefore agreement between annotators is much higher.

### v3

@@ V Conclusion
In this work, we evaluate whether CoT displays the same priors as simple ICL. We find, surprisingly, that larger language models do have reasoning priors that ossify their generated reasoning and posterior predictions, virtually disregarding the evidence in the form of reasoning and labels provided in the prompt. Given that the performance of LLMs is inferior to traditional methods in the tasks we study, this is unlikely to be an artifact of data memorization. We conclude that reasoning in the form of CoT is unable to overcome the pull of the priors in larger language models, leading to suboptimal performance on subjective tasks in the form of multilabel benchmarks, as opposed to other tasks where labels, and in turn reasoning, remain more consistent across datasets, the validity of a label can be asserted more formally and consistently, and therefore agreement is much higher.